OpenStudy (anonymous):

A shoe store has a cost of 5x + 20 dollars and a revenue of -x2 + 4x + 6 dollars for every pair of shoes sold. Find the company's profit equation if Profit = Revenue - Cost. plz help me

OpenStudy (gw2011):

P=(-x^2+4x+6)-(5x+20) P=-x^2+4x+6-5x-20 P=-x^2-x-14 This is the profit equation